# Towards scalable active steering protocols for genuinely entangled state manifolds

### v1

@@ Abstract
Numerical simulation results.— We now show simulation results for the above protocol maximizing the QFI. For the steering operator set (2), it is convenient to choose ${\bf s}_{n}=\frac{1}{\sqrt{2}}(1,0,1)^{T}$, see Eq. (3), but our results are robust under small rotations of this unit vector. Correspondingly, the GHZ states (1) are defined with respect to rotated states $|0\rangle\to|0^{\prime}\rangle$ and $|1\rangle\to|1^{\prime}\rangle$. Up to a normalization factor, $|0^{\prime}\rangle=|0\rangle+(\sqrt{2}-1)|1\rangle$ and $|1^{\prime}\rangle=(1-\sqrt{2})|0\rangle+|1\rangle$. In Fig. 2, we show the evolution of the time-dependent QFI. The main panel illustrates the average QFI $\overline{F_{Q}(t)}$ (the overbar indicates an average over many measurement trajectories) for several values of $N$, where we observe that the QFI comes closes to its maximum value $F_{Q}=N^{2}$ after $n_{t}\approx 200$ time steps. This number for $n_{t}$ is basically independent of $N$. In the inset of Fig. 2, for $N=5$, we illustrate the convergence behavior of the QFI both for individual measurement trajectories and for the average. Next we show that the target manifold (1) is reached to good accuracy. In Fig. 3(a), for $N=5$, we show a histogram of the phase $\phi=\mathrm{arg}\left(\langle 0^{\prime}0^{\prime}0^{\prime}\cdots|\Psi\rangle\langle\Psi|1^{\prime}1^{\prime}1^{\prime}\cdots\rangle\right)$ measured after 500 time steps. We find that the histogram is rather flat, implying that the quantum state trajectories uniformly explore the entire manifold even though they all start from the same initial state. Alternatively, as illustrated in Fig. 3(b), one may take an individual measurement trajectory and follow it over time. For our choice of steering operators (2), almost regular oscillations are observed, where again all possible values of $\phi$ are reachable over the course of time. By invoking a termination policy, one can then target a specific state with a predesignated value of $\phi$. We note that the oscillatory behavior appears to be linked to the specific sign choice of $J$ in Eq. (2). Finally, in Fig. 3(c), we show the purity [34, 35, 36],
@@
Figure 4: Scaling of the accuracy of the asymptotic value of the averaged QFI $\overline{F_{Q}}/N$ vs $N$ for $J\delta t=0.1$ (blue) and $J\delta t=0.2$ (red). Results have been averaged over $100$ measurement trajectories. Solid lines are guides to the eye only. The upper bound, $\overline{F_{Q}}/N=N$, is shown as black dashed line, the green dashed line indicates $\overline{F_{Q}}/N=0.9N$.
Figure 5: Scaling of the step number $n_{t}$ vs $N$ needed for reaching the averaged QFI value $\overline{F_{Q}}=0.9N^{2}$, shown for $J\delta t=0.1$ and $J\delta t=0.2$. Averages are over $100$ trajectories. Solid and dashed curves are guides to the eye only. We show results for nearest-neighbor detector qubit measurements (solid) as well as allowing for arbitrary detector pair measurements (dashed curves).
@@
In Fig. 5, we report the scaling of the number of steps $n_{t}$ needed for reaching a QFI of $\overline{F_{Q}}/N^{2}=0.9$, for the same values of $J\delta t$ as in Fig. 4. Here, we also compare to a situation where one allows for Bell pair measurements between arbitrary detector pairs (not only nearest neighbors), which tends to accelerate the protocol. We observe an even-odd effect, in particular for small $N$, which originates from the presence of an idle qubit for odd $N$ in our Bell pair measurement scheme, see Fig. 1. Remarkably, the required number of steps $n_{t}$ increases only very slowly (at most polynomially) with $N$, suggesting scalability of the active steering protocol for large $N$. This behavior is also corroborated by Fig. 2, where we show the $n_{t}$-dependence of $\overline{F_{Q}}$ for several (large) values of $N$.
@@
Discussion.—We have proposed an active steering protocol targeting the one-parameter manifold of $N$-qubit GHZ states (1) with genuine multipartite entanglement by means of weak Bell pair measurements and active feedback. In contrast to fidelity-based cost functions [6], by using the QFI as a cost function, our results suggest that the active steering protocol becomes scalable with increasing system size $N$. Although it is well known that GHZ states may be generated by projective measurements in shallow circuits [38], our results are relevant for at least two reasons: (i) We clarify similarities and differences between active and passive steering protocols for the case of a target state manifold. In passive steering [9, 39], measurement outcomes are discarded. Such protocols are basically equivalent to driven-dissipative systems [40, 41, 42, 43, 44] with engineered dissipation, see Refs. [45, 46, 47, 48, 49, 50, 11, 51, 52, 53, 54, 55, 56, 57, 58, 59] and corresponding experiments  [60, 61, 62, 63, 64, 65]. In contrast to the passive steering case, where the initial state uniquely determines the final state within the dark space forming the target manifold [66], we find that under active steering, the quantum state trajectory continues cycling through the target manifold. The initial state then plays no special role, and all phases $\phi$ in Eq. (1) are reached with equal probability as the protocol evolves. (ii) One may also target highly entangled non-stabilizer state manifolds, for which simpler routes along the lines of Ref. [38] are not available. Under quite general conditions, many highly entangled states, including Eq. (1), cannot be reached by driven-dissipative and/or passive steering [67], while they are accessible to active steering protocols [6]. This task may be achieved, e.g., by adding a non-stabilizerness quantifier term [68, 69] to the cost function, or by stopping the protocol at a suitable value of $F_{Q}$ before convergence has been reached. Active steering could thus allow for a powerful approach in this case. However, we leave a detailed exploration to future work.
Furthermore, for an experimental implementation, a modified protocol would be highly beneficial where state tracking is not required. It may indeed be possible to formulate such a protocol if weak measurements of all one- and two-body correlations present in the system state $|\Psi\rangle$ can be performed, see also Ref. [11]. In fact, the average cost function change $\overline{dF_{Q}}$ in Eq. (Towards scalable active steering protocols for genuinely entangled state manifolds) depends only on those correlations.
Finally, given that the manifold achieved by our protocol allows one to overcome classical limits in quantum phase-estimation and to attain the so-called Heisenberg limit [16, 17, 18, 19, 20], it would be interesting to further explore how to use active steering for entanglement-enhanced metrology.

### v2

@@ III Numerical simulation results
We now show simulation results for the above protocol maximizing the QFI. For the steering operator set (2), it is convenient to choose ${\bf s}_{n}=\frac{1}{\sqrt{2}}(1,0,1)^{T}$, see Eq. (3), but our results are robust under small rotations of this unit vector. Correspondingly, the GHZ states (1) are defined with respect to rotated states $|0\rangle\to|0^{\prime}\rangle$ and $|1\rangle\to|1^{\prime}\rangle$. Up to a normalization factor, $|0^{\prime}\rangle=|0\rangle+(\sqrt{2}-1)|1\rangle$ and $|1^{\prime}\rangle=(1-\sqrt{2})|0\rangle+|1\rangle$. In Fig. 2, we show the evolution of the time-dependent QFI. The main panel illustrates the average QFI $\overline{F_{Q}(t)}$ (the overbar indicates an average over many measurement trajectories) for several values of $N$, where we observe that the QFI comes close to its maximum value $F_{Q}=N^{2}$ after $n_{t}\approx 200$ time steps. This number for $n_{t}$ is basically independent of $N$. In the inset of Fig. 2, for $N=5$, we illustrate the convergence behavior of the QFI both for individual measurement trajectories and for the average.
Next, we show that the target manifold (1) is reached to good accuracy. In Fig. 3(a), for $N=5$, we show a histogram of the phase $\phi=\mathrm{arg}\left(\langle 0^{\prime}0^{\prime}0^{\prime}\cdots|\Psi\rangle\langle\Psi|1^{\prime}1^{\prime}1^{\prime}\cdots\rangle\right)$ measured after 500 time steps. We find that the histogram is rather flat, implying that the quantum state trajectories uniformly explore the entire manifold even though they all start from the same initial state. Alternatively, as illustrated in Fig. 3(b), one may take an individual measurement trajectory and follow it over time. For our choice of steering operators (2), almost regular oscillations are observed, where again all possible values of $\phi$ are reachable over the course of time. By invoking a termination policy, one can then target a specific state with a predesignated value of $\phi$. Finally, in Fig. 3(c), we show the purity [48, 49, 50],
@@
Figure 4: Scaling of the accuracy of the asymptotic value of the averaged QFI $\overline{F_{Q}}/N$ vs $N$ for $J\delta t=0.1$ (blue) and $J\delta t=0.2$ (red). Results have been averaged over at least $100$ measurement trajectories. (For $N\leq 10$, we used 10000 trajectories; for $N=11$ to $15$, we averaged over 5000, and for $N\in\{16,17\}$, over 1000 trajectories.) Solid lines are guides to the eye only. The upper bound, $\overline{F_{Q}}/N=N$, is shown as black dashed line, the green dashed line indicates $\overline{F_{Q}}/N=0.9N$.
Figure 5: Scaling of the step number $n_{t}$ vs $N$ needed for reaching the averaged QFI value $\overline{F_{Q}}=0.9N^{2}$, shown for $J\delta t=0.1$ and $J\delta t=0.2$. Averages are over at least $100$ trajectories. Thick solid and dashed curves connecting data points are guides to the eye only. The thin solid lines show numerical fits to the function $n_{t}^{*}(N)=A+B\ln(N)$ to the respective data, with fitting parameters $A$ and $B$. For instance, we obtain $A\simeq 104.6$ and $B\simeq 124$ for nearest-neighbor couplings with $J\delta t=0.1$ from such a fit. We show results for nearest-neighbor detector qubit measurements (solid) as well as allowing for arbitrary detector pair measurements (dashed curves), see Fig. 1.
@@
In Fig. 5, we report the scaling of the number of steps $n_{t}$ needed for reaching a QFI of $\overline{F_{Q}}/N^{2}=0.9$, for the same values of $J\delta t$ as in Fig. 4. Here, we also compare to a situation where one allows for Bell pair measurements between arbitrary detector pairs (not only nearest neighbors), see Fig. 1(c), which tends to accelerate the protocol. In this case, all non-overlapping pairs $(n,n^{\prime})$ with $n,n^{\prime}\in\{1,\dots,N\}$ can again be steered simultaneously, where both the pairings (and the idle qubit for odd $N$) are chosen from a uniform random distribution for each time step. Such a fully connected Bell measurement pairing scheme results in a faster convergence to the target state since more options for finding optimal feedback Hamiltonians can be explored.
@@
We observe an even-odd effect in Fig. 5, in particular for small $N$, which originates from the presence of the idle qubit for odd $N$ in our Bell pair measurement scheme. Remarkably, the required number of steps $n_{t}$ increases only very slowly with $N$, suggesting scalability of the active steering protocol for large $N$. The fitting curves shown as thin solid lines in Fig. 5 indicate a scaling ${n_{t}\sim{\cal O}(\ln N)}$. Such favorable scaling is also corroborated by Fig. 2, where we show the $n_{t}$-dependence of $\overline{F_{Q}}$ for several (large) values of $N$.
Figure 6: Average QFI $\overline{F_{Q}}$ vs number of time steps $n_{t}=t/\delta t$ for an active steering protocol with cost function $|F_{Q}-F_{Q}^{\ast}|$ and $F_{Q}^{\ast}$ in Eq. (13), targeting the Dicke states $|D_{k,N}\rangle$ in Eq. (12) for $k=N/2$ and various $N$, with $J\delta t=0.2$. Averages are over $5000\,(1000,500)$ trajectories for $N=14\,(16,18)$. Dotted lines indicate the respective $F_{Q}^{\ast}$ in Eq. (13). The inset shows $\overline{F_{Q}}$ vs $n_{t}$ (solid curve, averaged over $10^{4}$ trajectories) for $N=4$, together with three individual measurement trajectories (dashed curves).
In addition to targeting states with maximum QFI, our protocol also allows for the preparation of states associated with different (non-maximal) target values $F_{Q}^{*}$ of the QFI. In order to do so, we simply choose $|F_{Q}-F_{Q}^{\ast}|$ as cost function in the steering protocol. As an example for this approach, we here consider the case of Dicke states [19], which for qubits can be written as
$|D_{k,N}\rangle=\left(\begin{array}[]{c}N\\ k\end{array}\right)^{-\frac{1}{2}}\sum_{j}\mathcal{P}_{j}(|0\rangle^{\otimes N-k}|1\rangle^{\otimes k}).$ (12)
In Eq. (12), the sum runs over all possible permutations ${\cal P}_{j}$ of distributing $k$ excited qubits in an $N$-qubit system. For the state in Eq. (12), the QFI has the value [19]
$F_{Q}^{\ast}=\frac{N^{2}}{2}-2\left(\frac{N}{2}-k\right)^{2}+N.$ (13)
As for GHZ states, the Dicke states in Eq. (12) can in addition contain arbitrary phase differences between the corresponding basis states, with the same value of $F_{Q}^{\ast}$. In general, our protocol will therefore target an entire state manifold for a given QFI value.
In analogy to Fig. 2, Figure 6 shows numerical simulation results for the QFI-based preparation of Dicke states with target value $F_{Q}^{\ast}$ in Eq. (13) for $k=N/2$ and up to $N=18$ qubits. We observe that the target value of the QFI is reached after $n_{t}\approx 70$ time steps, independently of the value of $N$. The protocol thus converges significantly faster than for GHZ states. However, even though $\overline{F_{Q}}$ is close to the target value $F_{Q}^{*}$, individual trajectories now exhibit stronger fluctuations, see the inset of Fig. 6. We note in passing that by implementing a stoppage criterion in the protocol [6], i.e., by terminating the protocol once the QFI target value has approximately been reached, we expect that such fluctuations can be reduced. In any case, we conclude that the QFI-based protocol is also useful for generating other highly entangled state manifolds beyond GHZ states.
IV Discussion
@@
We have proposed an active steering protocol targeting the one-parameter manifold of $N$-qubit GHZ states (1) with genuine multipartite entanglement by means of weak Bell pair measurements and active feedback. In contrast to fidelity-based cost functions [6], by using the QFI as a cost function, our results suggest that the active steering protocol may become scalable with increasing system size $N$. Although it is well known that GHZ states may be generated by projective measurements in shallow circuits [52, 23, 24], our results are relevant for at least two reasons beyond those specified in Sec. I: (i) We clarify similarities and differences between active and passive steering protocols for the case of a target state manifold. In passive steering [9, 53], measurement outcomes are discarded. Such protocols are basically equivalent to driven-dissipative systems [54, 55, 56, 57, 58] with engineered dissipation, see Refs. [59, 60, 61, 62, 63, 64, 11, 65, 66, 67, 68, 69, 70, 71, 72, 73] and corresponding experiments  [74, 31, 75, 76, 77, 78]. In contrast to the passive steering case, where the initial state uniquely determines the final state within the dark space forming the target manifold [79], we find that under active steering, the quantum state trajectory continues cycling through the target manifold. The initial state then plays no special role, and all phases $\phi$ in Eq. (1) are reached with equal probability as the protocol evolves. (ii) One may also target highly entangled non-stabilizer state manifolds, for which simpler routes along the lines of Ref. [52] are not available. Under quite general conditions, many highly entangled states, including Eq. (1), cannot be reached by driven-dissipative and/or passive steering [80]. For small $N$, it has been demonstrated that they remain accessible to active steering protocols [6].
In fact, one can apply a slightly modified version of the present protocol, which has favorable scaling properties with system size, in order to prepare broad classes of other states, including the so-called $W$, Dicke, and cluster states [21]. This task may be achieved, for instance, by stopping the protocol at a suitable value of $F_{Q}$ before the convergence to the maximal QFI has been reached. Alternatively, one can select a target value $F_{Q}^{\ast}$ for the QFI associated with the desired family of target states, and then choose $|F_{Q}-F_{Q}^{\ast}|$ as cost function. For instance, the Dicke states $|D_{k,N}\rangle$ in Eq. (12) with associated QFI $F_{Q}^{\ast}$ in Eq. (13) have been studied for $k=N/2$ excitations above, see Fig. 6. Another possibility to reach non-GHZ states is to add a non-stabilizerness quantifier [81, 82] to the cost function.
For experiments, a modified implementation of our protocol where state tracking is not required would be highly beneficial. Without state tracking, one could both study larger system sizes and tolerate error channels with weak error rates. Such protocols can be formulated if weak measurements of all one- and two-body correlations present in the system state can be performed [11] since $\overline{dF_{Q}}$ in Eq. (II) depends only on those correlations. As a result, the protocol could be entirely based on the outcomes of weak measurements without any need for state tracking. In particular, our simulations only use the values of the one- and two-body correlations $R_{\mu_{i}}$ and $Q_{n,m}^{\alpha_{n},\alpha_{m}}$ defined after Eq. (8) in order to determine the optimal couplings $K$ entering the feedback Hamiltonian. Counting the number of one- and two-body correlators, we expect that the performance of such a protocol will scale $\sim N^{2}$. Noting that our numerical simulations are mainly limited by the computational demands of updating the SSE, which is not necessary anymore when switching to a protocol directly utilizing weak measurements of one- and two-body correlations, we expect that such a protocol can be applied to larger values of $N$. In addition, it allows for the presence of error channels and/or measurement inefficiencies. However, a detailed exploration of this interesting direction is beyond the scope of this paper.
Finally, given that the state manifolds that can be targeted by our protocol could allow one to overcome classical limits in quantum phase estimation and to attain the so-called Heisenberg limit [16, 17, 18, 20], it is a promising option to employ active steering protocols using the QFI for entanglement-enhanced metrology and sensing applications, see also Refs. [83, 84, 19, 85]. However, this remains feasible provided errors have a limited impact and if an optimal measurement, dependent on the state, can be performed. A comprehensive analysis of how errors affect state preparation, sensing, and measurement is left for future work.
Data availability. The data underlying the figures and the source code used for generating the simulations are available at [86, 87].